# MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks

### v2

@@ Abstract
Large Language Models (LLMs) have showcased impressive capabilities in handling straightforward programming tasks. However, their performance tends to falter when confronted with more challenging programming problems. We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions. To overcome this limitation, we present Modular-of-Thought Coder (MoTCoder). We introduce a pioneering framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 12.9% on APPS and 9.43% on CodeContests. Our codes are available at https://github.com/dvlab-research/MoTCoder.
@@
itemsep=0pt,partopsep=0pt,parsep=0pt,topsep=0pt
[…]
1 Introduction
It has been a longstanding objective in the field of artificial intelligence to develop systems capable of generating executable and functionally correct computer programs to address intricate problems (Manna & Waldinger, 1971). Recently, there has been considerable attention on Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022), which showcases remarkable success. These models leverage extensive pre-training and undergo further fine-tuning with detailed instruction data (Ouyang et al., 2022), leading to SOTA performance across diverse tasks. Initially designed for natural languages, these models have been expanded through the instruction fine-tuning for code modeling capabilities (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c). This extension has led to impressive performance in generating code from natural language problem descriptions (Li et al., 2023; Luo et al., 2023a; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021).
Figure 1: APPS test results by pass@k (%). Comparison of our MoTCoder with previous SOTAs.
Figure 2: The illustration of our MoTCoder training framework. We first perform the modular-of-thought instructions evolution, including the filtration of instructions that do not undergo successful evolution. Subsequently, we delve into the MoT instruction tuning and testing of the model.
Figure 3: To evolve conventional instructions into modular-of-thought instructions, our MoT instruction evolution approach guides the models through a two-step process. Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. The subsequent instruction then guides the model to implement these modules and integrate them into a cohesive final solution. This instruction is complemented by a one-shot example to further enhance understanding.
However, when confronted with highly intricate coding challenges, the current SOTA models still struggle to match the seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023), mainly due to their overly simplistic generation approach. Many previous LLMs produce the code solution as a single monolithic block, instead of breaking down of the task into logical sub-tasks. However, tackling highly complex problems with a single code module is practically unattainable. In contrast, adept developers devise modularized solutions, organizing them into high-level logical sub-modules. Subsequently, they complete and combine these components to efficiently enhance their final solutions.
In our Modular-of-Thought Coder (MoTCoder), our objective is to enhance the modularization capabilities of coding Large Language Models (LLMs) by introducing modular-of-thought prompting. This approach guides LLMs to break down their solutions into modular segments, where each segment represents an abstract function dedicated to a high-level logical sub-task. To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Evolution.
As depicted in Figure 2, our MoT code instruction evolution framework comprises two key stages: In the initial stage, we tailor the evolutionary prompt process. LLMs are instructed to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instruction guides the model in implementing these modules and eventually combining them into a comprehensive final solution. In the second stage, we fine-tune the LLM using our created code instruction-following training set, resulting in our MoTCoder model.
Figure 4: Illustration of modular code generation involves the model initially delineating the necessary sub-modules, each comprising a function header and a docstring detailing its intended purpose. Following this, the model proceeds to comprehensively implement each module in code and incorporates them as integral components of the finalized solution..
Our experiments reveal that MoTCoder significantly enhances LLM performance, establishing new SOTA results on challenging code tasks in APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a). Specifically, MoTCoder elevates the pass@1 performance by over 12.9% on APPS and 9.43% on CodeContests.
@@ 2.1 Large Language Models
General LLMs. In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks. Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 and GPT4 (Brown et al., 2020; OpenAI, 2023), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude. The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available. EleutherAI has contributed GPT-NeoX-20B (Black et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). Google has released UL2-20B (Tay et al., 2022). Tsinghua University has introduced GLM-130B (Zeng et al., 2022). Meta has released OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023).
Coding LLMs. Recent research has introduced a significant number of LLMs tailored for code-related tasks to address the challenges of code understanding and generation. Closed-source models include OpenAI’s Codex (Chen et al., 2021a) and Code-Davinci (Microsoft, 2023). Google has proposed PaLM-Coder (Chowdhery et al., 2022). These models excel on popular code completion benchmarks such as HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b). Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023). These models have demonstrated significant advancements in code-related tasks.
@@ 2.2 Instruction Fine-Tuning
General Instruction Tuning. In its early stages, the core aim of instruction fine-tuning was to amplify the cross-task generalization capabilities of Language Models (LMs). This was accomplished by subjecting LMs to fine-tuning using an extensive corpus of public Natural Language Processing (NLP) tasks. Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks. Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs. Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines. OpenAI has taken an alternative route by enlisting human annotators to contribute an extensive corpus of human instructions, encompassing diverse formats and a broad spectrum of task types. Building upon this dataset, OpenAI trained its GPT-3 (Brown et al., 2020) model to create InstructGPT (Ouyang et al., 2022), which better aligns with users’ inputs. This developmental trajectory has given rise to notable works such as ChatGPT. In the open-source realm, Alpaca (Taori et al., 2023) adopts the self-instruct method (Wang et al., 2022), leveraging ChatGPT to generate data for training. Vicuna (Chiang et al., 2023) utilizes user-shared conversations collected from ShareGPT.com to train its models. Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets. In contrast to these generalized instruction fine-tuning approaches, WizardCoder (Luo et al., 2023b) applies the Evol-Instruct method specifically in the domain of Code LLMs.
Chain-of-Thought Instruction Tuning. While large language models have demonstrated impressive capabilities in handling straightforward programming tasks, their performance tends to diminish when faced with more challenging programming problems. A notable limitation is observed in conventional models, where solutions are often generated as monolithic code blocks, limiting their effectiveness in addressing intricate questions. Drawing inspiration from the success of chain-of-thought (CoT) prompting in enhancing performance in reasoning-based tasks (Zhou et al., 2023; Wei et al., 2022b; Kojima et al., 2022; Le et al., 2023), we introduce a pioneering framework for Modular-of-Thought (MoT) instruction tuning. MoT is designed to facilitate the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
@@ 3 Methods
The depiction of our MoTCoder training framework is presented in  Figure 2. In the initial stage, we introduce the modular-of-thought instructions evolution and evolution examination in  Section 3.1. Following this, we elaborate on the MoT instruction tuning and testing of the model in  Section 3.2.
3.1 Modular-of-Thought Instruction Evolution
In this section, we propose the modular-of-thought instruct evolution approach. We initiate the evolution process with a given initial instruction dataset $D^{(0)}={(I_{n}^{(0)},\hat{O}_{n}^{(0)})}_{1\leq n\leq N}$, where $I_{n}^{(0)}$ represents the $n$-th instruction in $D^{(0)}$, $O_{n}^{(0)}$ is the corresponding output, and $N$ is the number of samples in $D^{(0)}$. In each evolution step, we upgrade all $I^{(k)}$ in $D^{(k)}$ to $I^{(k+1)}$ by employing an LLM instruction evolution prompt. Subsequently, we use the LLM to generate corresponding outputs $\hat{O}^{(k+1)}$ for the newly evolved $I^{(k+1)}$. Thus, we obtain an evolved instruction dataset $\{D^{(0)}\dots D^{(k+1)}\}$. By iteratively conducting $K$ evolutions, we sequentially acquire the final MoT instruction dataset $\{D^{(0)}\dots D^{(K)}\}$.
Our process for the evolution of modular-of-thought instructions comprises two key stages: MoT instructions evolution, and evolution examination, involving the filtration of instructions that do not successfully evolve. We commence by populating the instruction pool with the provided initial instruction dataset $D^{(0)}$. In each subsequent evolution epoch, normal instructions from the preceding epoch are retrieved from the pool. Subsequently, we employ the instruction evolver to advance each retrieved instruction and the instruction eliminator to evaluate whether the evolution is unsuccessful. Instructions that evolve successfully are integrated into the final pool, while unsuccessful ones are returned as they are, with the anticipation of achieving successful enhancement in the following evolution epoch.
Normal Instruction.
@@ 3.1 Modular-of-Thought Instruction Evolution
$\hat{o}_{t}\sim p_{\theta}(.|\hat{o}_{1:t-1},I),$ (1)
where $I$ represents the input instruction and $\hat{o}_{t}$ is the $t$-th of the flattened output sequence.
Modular-of-Thought Instruction.
Referring to the recent works in code revision (Le et al., 2023), our proposed methodology aims to evolute normal instructions into a sequential code generation process by leading the models through a two-step procedure.
@@
    Final solution. The subsequent instruction guides the model to implement these modules and eventually combine them into a comprehensive final solution.
@@
Modular-of-Thought Instruction Illustration.
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 3 and further detailed in Figure 4. The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
(a) Modular Prompt
(b) Normal Prompt
Figure 5: Example of modular and normal code generation prompts. The modular instructions provide sub-modules including their function headers and docstrings outlining their intended usage. The final resolution executes these modules and then incorporates them into a comprehensive solution.
Evolution Examination.
We categorize the following scenarios as instances of instruction evolution failure:
@@
    The evolved instruction deviates from the modular-of-thought generation strategy, failing to follow the approach of initially generating sub-modules and then generating the main code.
@@
    During the first phase, namely the sub-module generation stage, either no sub-modules are generated, or global code is produced.
@@
    In the second phase, or the main code generation stage, there is either a lack of main code generation or the production of multiple main code segments.
Following the evolution process, we created a modular-of-thought instruction dataset comprising around 24k instructions. The dataset exhibits a distribution ratio of 45:55 between MoT and normal instruction data.
3.2 Modular-of-Thought Training and Inference
We present the MoT prompt training methodology through the introduction of a modular training prompt, as illustrated in Figure 5a. The instructions first specify sub-modules with their function headers and docstrings detailing their intended usage. The ultimate solution involves implementing these modules and subsequently integrating them into a final solution. The {input} in prompts is the problem description and the model is trained to complete the {response} in prompts. We conducted instruction tuning using the base model WizardCoder (Luo et al., 2023b) on our MoT instruction dataset for three epochs. This process took approximately 28 hours using 8 A800 GPUs.
Table 1: Experiments of modular ability. We conducted inference using both normal and modular prompts, as demonstrated in Figure 5 and show APPS test results by pass@1 (%). Comparable approaches include StarCoder (Li et al., 2023) and WizardCoder (Luo et al., 2023b) with the same size of 15B.
Model Size Prompt Introductory Interview Competition All
StarCoder 15B Normal 7.25 6.89 4.08 6.40
Modular 0.59 2.87 1.00 2.04
WizardCoder 15B Normal 26.04 4.21 0.81 7.90
Modular 0.39 0.46 0.66 0.81
Normal 33.80 19.70 11.09 20.80
MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 2: Modular-of-thought instruction tuning ablation experiments. We compare the one-step and two-step modular-of-thought instruction tuning strategies. Models are inferred using both normal and modular prompts. APPs test results by pass@1 (%) are shown.
Model Size Prompt Introductory Interview Competition All
Normal 30.76 17.16 11.16 18.68
1-Step MoTCoder 15B Modular 31.69 17.37 11.47 19.05
Normal 33.80 19.70 11.09 20.80
2-Step MoTCoder 15B Modular 35.86 18.79 11.82 20.81
Table 3: Training Datasets. We compare MoTCoder training on only modular dataset and the mixture of modular and normal dataset. APPs test results by pass@1 (%) are shown.
Model Introductory Interview Competition All
Baseline 7.25 6.89 4.08 6.40
Modular Only 31.34 15.16 10.25 17.42
Mixture 33.80 19.70 11.09 20.80
Table 4: APPS test results by pass@k (%). Competitive open-source approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). We also present results from closed-source models, namely text-davinci, code-davinci, and GPT3.5.
Model Size Pass@ Introductory Interview Competition All
CodeT5 770M 1 6.60 1.03 0.30 2.00
GPT-Neo 2.7B 1 14.68 9.85 6.54 10.15
5 19.89 13.19 9.90 13.87
GPT-2 0.1B 1 5.64 6.93 4.37 6.16
5 13.81 10.97 7.03 10.75
1.5B 1 7.40 9.11 5.05 7.96
5 16.86 13.84 9.01 13.48
GPT-3 175B 1 0.57 0.65 0.21 0.55
StarCoder 15B 1 7.25 6.89 4.08 6.40
WizardCoder 15B 1 26.04 4.21 0.81 7.90
MoTCoder 15B 1 33.80 19.70 11.09 20.80
text-davinci-002 - 1 - - - 7.48
code-davinci-002 - 1 29.30 6.40 2.50 10.20
GPT3.5 - 1 48.00 19.42 5.42 22.33
Table 5: APPS test results with code-revision by pass@1 (%). Revision indicates the rounds of revision after the first round of direct generation. Baselines include CodeT5 (Wang et al., 2021), text-davinci and WizardCoder (Luo et al., 2023b) and code-revision methods contain Self-edit (Zhang et al., 2023), CodeRL (Le et al., 2022), Self-repair (Olausson et al., 2023), and CodeChain (Le et al., 2023).
Model Size Revision Introductory Interview Competition All
CodeT5 770M - 6.60 1.03 0.30 2.00
CodeRL+CodeT5 770M 5 7.08 1.86 0.75 2.69
text-davinci-002 - - - - - 7.48
Self-edit+text-davinci-002 - 5 - - - 7.94
WizardCoder 15B - 26.04 4.21 0.81 7.90
CodeChain+WizardCoder 15B 5 26.29 7.49 3.75 10.50
MoTCoder 15B - 33.80 19.70 11.09 20.80

### v3

# MoTCoder: Elevating Large Language Models with Module-of-Thought

@@ Abstract
Large Language Models (LLMs) have showcased impressive capabilities in handling straightforward programming tasks. However, their performance tends to falter when confronted with more challenging programming problems. We observe that conventional models often generate solutions as monolithic code blocks, restricting their effectiveness in tackling intricate questions. To overcome this limitation, we present Module-of-Thought Coder (MoTCoder). We introduce a framework for MoT instruction tuning, designed to promote the decomposition of tasks into logical sub-tasks and sub-modules. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly improves both the modularity and correctness of the generated solutions, leading to substantial relative pass@1 improvements of 60% on APPS and 19% on CodeContests. Our codes are available at https://github.com/dvlab-research/MoTCoder.
1 Introduction
Developing systems that can generate executable and functionally correct computer programs has long been sought after in artificial intelligence (Manna & Waldinger, 1971). Recently, Large Language Models (LLMs) (Brown et al., 2020; OpenAI, 2023a; Chowdhery et al., 2022; Anil et al., 2023; Hoffmann et al., 2022; Rae et al., 2021; Zeng et al., 2022; Touvron et al., 2023; Zhang et al., 2022) have showcased remarkable success in many problem domains beyond natural language processing, and is poised as a promising approach to tackle code modeling and generation (as well as other coding related tasks) (Rozière et al., 2023; Black et al., 2021; Chen et al., 2021c). Through instruction fine-tuning (Li et al., 2023b; Luo et al., 2023b; Wang et al., 2023a; Li et al., 2022b; Nijkamp et al., 2023; Zheng et al., 2023; Fried et al., 2022; Chen et al., 2021a; Wang et al., 2021), LLMs have achieved impressive performance on code generation benchmarks like HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021).
Yet, when confronted with intricate coding problems such as APPS (Hendrycks et al., 2021b) and CodeContests  (Li et al., 2022a), SOTA models struggle to match seasoned developers (Hendrycks et al., 2021b; Li et al., 2022a; Shinn et al., 2023). The main culprit is their overly simplistic generation approach. Current models produce the code solution as a single monolithic block in an attempt to solve the problem in one shot. While feasible for simple tasks, this approach becomes increasingly inefficient to tackle a complex task which naturally entails solving multiple sub-tasks. In contrast, adept developers often devise modularized solutions by breaking down the original problem into more approachable components that can be individually and more efficiently solved.
Following this intuition, recent works (Jiang et al., 2023; Le et al., 2023) propose iterative code inference for code generation. They, however, come with added inference costs and offer only marginal performance gains. In this work, we propose to more efficiently improve the modularization capability of coding LLMs using Module-of-Thought (MoT) instruction fine-tuning. Our approach guides LLMs to break down their solution into modular components, each representing an abstract function dedicated to a logical sub-task. To train the model to adhere to the MoT prompt, we generate instructional data using a process termed MoT Code Instruction Transformation. In this process, LLMs are instructed to outline necessary modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instruction guides the model to implement these modules and eventually combine them into the final solution. After that, we fine-tune the LLM with our MoT instruction fune-tuning, resulting in our MoTCoder model.
In summary, our contributions are threefold:
  1. 1.
    We propose a 2-step Module-of-Thought Code Instruction Transformation approach that instructs LLMs to generate modularized code solutions, where each module is to address a logical sub-task and accompanied by standardized annotations.
  2. 2.
    We develop Module-of-Thought Coder (MoTCoder), a new model that enhances the modularization capabilities of LLMs with MoT Instruction Tuning. It breaks down complex problems into manageable, logically coherent sub-modules, improving both comprehensibility and maintainability of the generated code.
  3. 3.
    Our experiments demonstrate that MoTCoder establishes new SOTA results on challenging code generation benchmarks such as APPS and CodeContests. Specifically, MoTCoder improves the pass@1 performance by significant margins, exemplified by improvements over existing metrics by 60% on APPS and 19% on CodeContests, as illustrated in Section 1.
@@ 1 Introduction
figurePass@1 results on CodeContests (left) and APPS (right). Comparison of our MoTCoder with previous SOTAs.
@@ 2.1 Large Language Models
General LLMs. In recent times, Large Language Models (LLMs) have exhibited remarkable prowess across a wide array of tasks. Leading technology companies have made significant advancements in developing highly proficient closed-source LLMs, including OpenAI’s GPT3 (Brown et al., 2020) and GPT4 (OpenAI, 2023a), Google’s PaLM (Chowdhery et al., 2022; Anil et al., 2023), Bard [^1] https://bard.google.com/, DeepMind’s Chinchilla (Hoffmann et al., 2022), and Gopher (Rae et al., 2021), as well as Anthropic’s Claude [^2] https://www.anthropic.com/index/introducing-claude. The AI community has also observed the release of several open-source LLMs, where model weights are made publicly available. EleutherAI has contributed GPT-NeoX-20B (Black et al., 2022) and GPT-J-6B (Wang & Komatsuzaki, 2021). Google has released UL2-20B (Tay et al., 2022). Tsinghua University has introduced GLM-130B (Zeng et al., 2022). Meta has released OPT (Zhang et al., 2022) and LLaMA (Touvron et al., 2023).
Coding LLMs. Recent research has introduced a significant number of LLMs tailored for code-related tasks to address the challenges of code understanding and generation. Closed-source models include OpenAI’s Codex (Chen et al., 2021a) and Code-Davinci (Microsoft, 2023). Google has proposed PaLM-Coder (Chowdhery et al., 2022). These models excel on popular code completion benchmarks such as HumanEval (Chen et al., 2021b) and MBPP (Austin et al., 2021). On the open-source front, Salesforce has introduced CodeGen (Nijkamp et al., 2023), CodeT5 (Wang et al., 2021), and CodeT5+ (Wang et al., 2023b). Tsinghua University has contributed CodeGeeX (Zheng et al., 2023), and the BigCode Project has developed StarCoder (Li et al., 2023b). These models have demonstrated significant advancements in code-related tasks.
@@ 2.2 Instruction Fine-Tuning
General Instruction Tuning. In its early stages, the core aim of instruction fine-tuning was to amplify the cross-task generalization capabilities of Language Models (LMs). This was accomplished by subjecting LMs to fine-tuning using an extensive corpus of public Natural Language Processing (NLP) tasks. Pioneering this approach, T5 (Raffel et al., 2020) underwent training on a diverse set of supervised text-to-text tasks. Subsequent endeavors like FLAN (Wei et al., 2022a), ExT5 (Aribandi et al., 2022), T0 (Sanh et al., 2022a), and UnifiedQA (Khashabi et al., 2020) broadened the spectrum of tasks, fortifying the overall generalization capability of LMs. Noteworthy contributions from ZeroPrompt (Xu et al., 2022) and FLAN-T5 (Chung et al., 2022) pushed boundaries by incorporating thousands of tasks into their training pipelines. OpenAI has taken an alternative route by enlisting human annotators to contribute an extensive corpus of human instructions, encompassing diverse formats and a broad spectrum of task types. Building upon this dataset, OpenAI trained its GPT-3 (Brown et al., 2020) model to create InstructGPT (Ouyang et al., 2022), which better aligns with users’ inputs. This developmental trajectory has given rise to notable works such as ChatGPT. In the open-source realm, Alpaca (Taori et al., 2023) adopts the self-instruct method (Wang et al., 2022), leveraging ChatGPT to generate data for training. Vicuna (Chiang et al., 2023) utilizes user-shared conversations collected from ShareGPT.com to train its models. Introducing the Evol-Instruct method, WizardLM (Xu et al., 2023) involves evolving existing instruction data to generate more intricate and diverse datasets.
Chain-of-Thought Instruction Tuning. In contrast to these generalized approaches, Recent works (Luo et al., 2023c; Yue et al., 2023; Chen et al., 2022b; Gunasekar et al., 2023; Haluptzok et al., 2023) apply instruction tuning method in common-sense reasoning (West et al., 2022), text-summarization (Sclar et al., 2022), mathematical reasoning  (Luo et al., 2023a; Yue et al., 2023), tool use (Patil et al., 2023), coding (Luo et al., 2023b), and general-purpose reasoning (Li et al., 2023a; Zelikman et al., 2022). Among them,  Yue et al. (2023) curates a diverse corpus of mathematics problems with chain-of-thought or program-of-thought (Chen et al., 2022b) annotations for mathematical reasoning analogous to our plans. Gunasekar et al. (2023) proposed pre-training models on programming textbooks generated synthetically from GPT3.5. Haluptzok et al. (2023) similarly generates programming puzzles and corresponding solutions from language models.
2.3 Prompting Techniques
A prompt technique known as Chain of Thought (CoT) (Wei et al., 2022b) generated intermediate reasoning steps before the answer to address language reasoning tasks. Inspired by CoT, a series of prompting works has been proposed. Least-to-most prompting (Zhou et al., 2022) reduces a complex problem into a series of sub-problems and then solves the sub-problems in order, adding the answer to the previous sub-problems to the prompt each time solving begins. PAL (Gao et al., 2022) and PoT (Chen et al., 2022a) is proposed to generating code as the intermediate reasoning steps, delegating solving to the compiler, thus improving solution accuracy. Similar methods are proposed for simple mathematical (Lewkowycz et al., 2022; Wu et al., 2022), commonsense (Sanh et al., 2022b; Madaan et al., 2022), symbolic reasoning (Yao et al., 2023) and code generation problems (Jiang et al., 2023; Le et al., 2023). However, these works can only plan code during generation. In comparison, our approach introduces a guided module-of-thought framework during training, making it more intrinsic. Our investigations reveal that, through the cultivation and utilization of sub-modules, MoTCoder significantly enhances both the modularity and correctness of the generated solutions.
Figure 1: Left: To transform conventional instructions into module-of-thought instructions, our MoT instruction transformation approach guides the models through a two-step process. Initially, the models are instructed to outline the necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. The subsequent instruction then guides the model to implement these sub-modules and integrate them into a cohesive final solution. This instruction is complemented by a one-shot example to further enhance understanding. Right: Examples of instruction, modularized code and solution.
@@ 3 Methods
In this section, we first detail the module-of-thought instruction transformation in Section 3.1 and then introduce the module-of-thought instruction tuning strategy for our MoTCoder in Section 3.2.
3.1 Module-of-Thought Instruction Transformation
In this section, we first introduce the normal instruction, followed by our module-of-thought instruction and its assessment.
Figure 2: Illustration of our 2-step Module-of-Thought Instruction Transformation: Plain code utilizes a single module to directly generate code. In contrast, our Module-of-Thought Instruction Transformation first instructs LLMs to outline necessary sub-modules, generating only their function headers and docstrings that describe their intended usage. Subsequently, the instructions guide the model in implementing these sub-modules and eventually combining them into a comprehensive final solution.
3.1.1 Normal Instruction
@@
$o_{t}\sim p_{\theta}(.|o_{1:t-1},I),$ (1)
where $I$ represents the input instruction and $o_{t}$ is the $t$-th token of the flattened output sequence.
3.1.2 Module-of-Thought Instruction
Considering programming encompasses both detailed coding abilities and broader logical or strategic planning capabilities, we propose a transformation pipeline that enhances code clarity and structure targeting detailed coding proficiencies and integrating descriptions based on natural language for the broader strategic planning capabilities.
According to previous researches (Jain et al., 2023), straightforward and unambiguous, detailed instructions enhance the model’s efficacy and precision in executing the desired tasks. Hence, for intricate data transformation tasks, decomposing the task into simpler, sequential steps yields better outcomes. Therefore, our proposed methodology aims to transform normal instructions into a sequential code generation process by leading the models through a two-step procedure, as illustrated in Figure 2 .
@@ 3.1.2 Module-of-Thought Instruction
    Final solution. The subsequent instruction guides the model to implement these sub-modules and eventually combine them into a comprehensive final solution.
@@
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 1. The instruction encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
3.1.3 Instruction Assessment
Throughout these transformations, we introduce guidelines for reviewing our transformed module-of-thought code. We identify the situations below as markers of instruction refinement failure:
@@ 3.1.3 Instruction Assessment
    The refined instruction diverges from the module-of-thought generation strategy, not adhering to the protocol of initial sub-module creation followed by the main code development.
@@
    At the sub-module creation phase, if no sub-modules are formed or if overarching code is developed instead.
@@
    During the main code development phase, the absence of main code creation or the emergence of multiple main code blocks indicates a problem.
  4. 4.
    The presence of test cases in the dataset that the transformed program fails to pass. This criterion ensures the transformed programs preserve functional equivalence with the original codes.
3.2 Module-of-Thought Instruction Tuning
In this section, we first introduce our module-of-thought dataset, followed by our MoT instruction tuning details.
3.2.1 MoT Dataset
We use data sourced from a small part of the training set of APPS (Hendrycks et al., 2021b) and CodeContests (Li et al., 2022a) as our initial set. For each problem, we take at most one answer, whereas the original dataset may contain dozens of answers. We then use GPT3.5 to generate transformed instructions, with the instruction transformation prompt as shown in Figure 1. The generated module-of-thought instruction dataset comprises around 24k instructions. We merge the docstrings that describe the intended usage and input-output variables type of the functions with their implementation. Some examples from the MoT dataset are in Appendix B.
3.2.2 MoT Instruction Tuning
Our MoTCoder underwent instruction tuning utilizing the SOTA coding model Deepseek-Cdoer-Base (Balog et al., 2016) across three epochs on our proposed MoT instruction dataset. The model’s maximum input length is 2048 tokens. We used a training batch size of 16 and an evaluation batch size of 1 per device, with gradient accumulation steps set to 4. The learning rate was initialized at $2\times 10^{-5}$, with a warmup period of 10 steps to gradually adapt the learning rate, following a cosine learning rate scheduler for optimization. Additionally, we enabled gradient checkpointing, BF16 precision and more deepseed settings to enhance training efficiency and reduce memory consumption: Our model was configured with the AdamW optimizer and utilized the WarmupLR scheduler to manage the learning rate adjustments effectively. To optimize memory and compute resources, we employed a third-stage zero optimization setting and enable communication overlap, contiguous gradients, and large sub group size of $1\times 10^{9}$ to streamline the training process. In addition, the maximum live parameters, maximum reuse distance, and the parameter settings for gathering 16-bit weights during model saving were all set to $1\times 10^{9}$. This process took approximately 28 hours using 8 A800 GPUs.